OpenStudy (anonymous):

Simplify quantity x squared plus 8 times x all over x squared. x plus 8 all over x squared x plus 8 all over x 8x 8

OpenStudy (jhannybean):

quantity x squared = x^2 8 times x = 8x quantity plus 8 times x = x^2 + 8x all over x squared = / x^2

OpenStudy (anonymous):

Then you have to simplify it o;

OpenStudy (jhannybean):

\[\large \frac{x^2+8x}{x^2} \]

OpenStudy (jhannybean):

whats the GCF of the numerator?

OpenStudy (anonymous):

x?

OpenStudy (jhannybean):

yes :D can you factor that out? factor out the smallest one. What do you get?

OpenStudy (anonymous):

x + 8 ----- x

OpenStudy (jhannybean):

\[\large \frac{x^2+8x}{x^2}= \frac{x(x+8)}{x^2}\]

OpenStudy (jhannybean):

You got it.

OpenStudy (anonymous):

^.^ thanks c:

OpenStudy (jhannybean):

Good job :D

OpenStudy (anonymous):

Hey umm how do you write equations like that?o;

OpenStudy (jhannybean):

the little "equation" button below the message box ---vv

OpenStudy (anonymous):

Can you tell me what to press on the equation tab and such i want to practice so i know for the future ;p like we say that equation in the problem

OpenStudy (anonymous):

\[\frac{ x^2 + 8x }{ x^2 }\]